OpenStudy (anonymous):

p(t)=t+3/2; find p(1/3)

OpenStudy (anonymous):

as in the equation we have variable t z given... n we have to take t=1/3

OpenStudy (anonymous):

...

OpenStudy (anonymous):

\[p(1/3)= 1/3 + 3/2 => p(1/3) =2+9 / 11\]

OpenStudy (anonymous):

p(1/3)= 2+9/ 6

OpenStudy (anonymous):

11 is wrng 6 is rite

OpenStudy (anonymous):

so p(t)= 11/6

OpenStudy (anonymous):

thank you